# FedHCDR: Federated Cross-Domain Recommendation with Hypergraph Signal Decoupling

### v2

@@ 2.3 High/Low-Pass Hypergraph Filter
According to the graph signal processing theory [5, 6], given a signal $\boldsymbol{x}\in\mathbb{R}^{|\mathcal{V}|}$ on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a filter $\boldsymbol{h}$, the graph convolution/filtering operator in the spectral (frequency) domain (denoted as *$*_{\mathcal{G}}$) is defined as
@@ 2.4 Local-Global Bi-Directional Transfer Algorithm
In this section, we introduce our proposed local-global bi-directional transfer algorithm. As shown in algorithm 1, in each round, the server sends the current global low-pass filter parameters $\boldsymbol{\phi}^{\mathrm{u},(t)}$ and global domain-shared user representations $\mathbf{U}^{\mathrm{g},(t)}$ to clients. In the local training stage, client $k$ updates its local high-pass hypergraph filter $\text{HHF}_{k}$ and low-pass hypergraph filter $\text{HLF}_{k}$ in an alternating manner, to perform knowledge transfer of global $\rightarrow$ local and local $\rightarrow$ global, as shown in lines $5-10$ of the pseudocode. This process allows for a trade-off between global and local models.
@@ 2.5 Hypergraph Contrastive Loss
To better obtain domain-invariant user relationship information, when updating the low-pass user hypergraph filter, we do perturbation to the user hypergraph by randomly dropping a portion of edges. First, we sample a edge random masking matrix $\mathbf{M}_{k}^{\text{u}}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{U}|}$, where $\left(\mathbf{M}_{k}\right)_{ij}\sim\text{Bernoulli}(1-p^{\text{d}})$ indicates whether to drop the edge between nodes $i$ and $j$. Here $p^{\text{d}}$ is the probability of each edge being dropped. Then the perturbed user hypergraph adjacency matrix can be computed as:
$\widetilde{\mathbf{A}}^{\text{u}}_{k}=\mathbf{A}^{\text{u}}_{k}\odot\mathbf{M}^{\text{u}}_{k},$ (19)
@@ 5.0.1 Acknowledgements
This work is supported by the National Natural Science Foundation of China (62227808), and Shenzhen Science and Technology Program (Grant No.ZDSYS20210623091809029).

### v3

@@ 2.3 High/Low-Pass Hypergraph Filter
According to the graph signal processing theory [5, 6], given a signal $\boldsymbol{x}\in\mathbb{R}^{|\mathcal{V}|}$ on a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a filter $\boldsymbol{h}$, the graph convolution/filtering operator in the spectral (frequency) domain (denoted as $*_{\mathcal{G}}$) is defined as
@@ 2.4 Local-Global Bi-Directional Transfer Algorithm
In this section, we introduce our proposed local-global bi-directional transfer algorithm. As shown in algorithm 1, in each round, the server sends the current global low-pass filter parameters $\boldsymbol{\phi}^{\mathrm{u},(t)}$ and global domain-shared user representations $\mathbf{U}^{\mathrm{g},(t)}$ to clients. In the local training stage, client $k$ updates its local high-pass hypergraph filter $\text{HHF}_{k}$ and low-pass hypergraph filter $\text{LHF}_{k}$ in an alternating manner, to perform knowledge transfer of global $\rightarrow$ local and local $\rightarrow$ global, as shown in lines $5-10$ of the pseudocode. This process allows for a trade-off between global and local models.
@@ 2.5 Hypergraph Contrastive Loss
To better obtain domain-invariant user relationship information, when updating the low-pass user hypergraph filter, we do perturbation to the user hypergraph by randomly dropping a portion of edges. First, we sample a edge random masking matrix $\mathbf{B}_{k}^{\text{u}}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{U}|}$, where $\left(\mathbf{B}_{k}\right)_{ij}\sim\text{Bernoulli}(1-p^{\text{d}})$ indicates whether to drop the edge between nodes $i$ and $j$. Here $p^{\text{d}}$ is the probability of each edge being dropped. Then the perturbed user hypergraph adjacency matrix can be computed as:
$\widetilde{\mathbf{A}}^{\text{u}}_{k}=\mathbf{A}^{\text{u}}_{k}\odot\mathbf{B}^{\text{u}}_{k},$ (19)
@@ 5.0.1 Acknowledgements
This work is supported by Guangdong Major Project of Basic and Applied Basic Researche (No. 2019B03030200245), the National Natural Science Foundation of China (62227808), and Shenzhen Science and Technology Program (Grant No.ZDSYS20210623091809029).